OpenStudy (anonymous):

Solve for “?” by finding a pattern that utilizes every number within each arrangement. The third arrangement will have the same pattern as the first two arrangements. 3 9 8 6 =14 10 2 7 3=16 11 8 17 4 =? "? can be 19,30, 22 or 32"

OpenStudy (anonymous):

3+9+8-6 =14 10+2+7-3 = 16 so 11+8+17-4 = ?